What is the Oxidation State of Sulphur in 


H2S2O8   ??

Grade:11

5 Answers

bibhash jha
14 years ago
+6. This is known as peroxy di sulphuric acid(marshals acid) where peroxylinkage is there ,so six oxygen normal with -2 charge while two with -1 chrge due to peroxy linkage so sulphur oxidation number is +6 .HO-S-O-O-S-OH both sulphur attach with two cordinate oxygen atom each.
Pritish Kamath
14 years ago

0=2*1+2x+ (-2)*8

where x is the oxidation state

therefore the oxidation state is 14

correct me i i am wrong somewhere

Regards

Ramesh V
14 years ago

O.S for H is +1

O.S for O is -2

let O.S for S be x , then

2+2x-16=0

x=+7  is O.S of sulphur

 

 

anandpal shekhawat
14 years ago
Hi, in this the two S atoms peroxy linkage between them as shown; (-S-O-O-S-) & each S has 2 double bonded O atoms & 1 -OH group. Hence, O.N. of S =+6.
Apoorv Jain
13 years ago

OXIDATION STATE OF SULPHUR IS 7